# Semantic Interaction Information mediates compositional generalization in latent space

### 28 Mar 2026 v1

@@ A.4 Learning, Dynamics and Computation
While there exists interaction information between observations $o^{i}$, interactions $z^{i}$, and realizations $r$- $I(o^{i};z^{i};r)=-I(r;z^{i}\mid o^{i})$, crucially, because different realizations of the same variable are mutually exclusive, the likelihood of a specific $r=r^{\prime}$ is perfectly isolated from all other potential realizations and $P_{Z}(r=r^{\prime}\mid\mathbf{o})$ may be updated independently. Next, consider a context of two latent variables ($C=2$). Here the interactions $Z^{i}=(z^{i}_{1},z^{i}_{2})$ map to a joint space via the bilinear tensor product $f(z^{i}_{12})\otimes f(z^{i}_{21})\in\mathbb{R}^{R\times R}$, forming an $R\times R$ matrix of belief updates. The bilinear construction generates synergy— a phenomenon where an informative combination of variables loses predictive power if the variables are considered in isolation, i.e. the whole is greater than the sum of its parts [73, 6]. We can see this because evaluating the likelihood of $o^{i}=1$ given a realization $r_{1}=r^{\prime}$ now involves projecting all $\mathbf{e}_{r_{2}}$ indicator vectors onto a shared basis:
$P_{Z^{i}}(o^{i}=1\mid r_{1}=r^{\prime},r_{2})=\mathbf{e}_{r^{\prime}}^{T}\sum_{r_{2}}\left(f(z^{i}_{1})f(z^{i}_{2})^{T}\right)\mathbf{e}_{r_{2}},$
revealing the entanglement between latent variables.
@@ A.5 Representation Classification Chains and Thousand Brains Theory
The Thousand Brains Theory of intelligence [33] postulates that individual cortical columns autonomously model through compositions of compositions. Tasked specifically with learning useful relationships, Representation Classification Chains can be viewed as the potential microcircuit of such a system. This architecture-agnostic circuit generates abstractions $\mathbf{r}$, to explain inputs $\mathbf{o}$, in a manner that enables optimal control of inputs given their internal value, without needing to interact with the environment. These chains incorporate the relational information of embeddings, decision-making of Classifiers, latent representation of generative models and embodiment of a Controller. Importantly, these modules are not combined arbitrarily, but have a principled grounding in the task of compositional generalization.
@@ A.6 Environment Hyperparameters
  • $|\mathcal{S}|$ (Total latent states): 500
@@ A.7 Additional Model Hyperparameters
  • Hidden dimensionality of all networks: 1000
@@
  • RCC Entropy bonus ($\beta_{1}$): $0.1$.
@@
  • RCC Entropy bonus ($\beta_{1}$): $0.01$.
  • Controller Learning rate: 0.001
  • Controller Training episodes: 10000
@@
  • Controller Entropy bonus ($\beta_{2}$): $0.999^{E}$
@@
When compositional embeddings are learned, training reliability was improved when the learnable embedding estimates were over-parameterized. The active state embedding estimates were passed through learnable linear layers $f_{K}(\cdot),f_{Q}(\cdot)\in\mathbb{R}^{d_{\hat{\mathcal{E}}}\times d_{\mathcal{E}}}$, to project them down from $d_{\hat{\mathcal{E}}}=N$ to dimensionality $d_{\mathcal{E}}$. Estimated interactions were computed from the projections:
@@
$\mathcal{L}_{\text{reg}}=\frac{1}{C\times d_{o}}\sum_{c,o}(\|f_{K}(\hat{\mathbf{k}}^{i}_{c})\|_{2}-1)^{2}+(\|f_{Q}(\hat{\mathbf{q}}_{c}^{i})||_{2}-1)^{2}$
@@
    LSTM: $h_{t}^{stm},h_{t}^{ltm}=\text{LSTM}(h_{t}^{in},(h_{t-1}^{stm},h_{t-1}^{ltm}))$ processes the $T\times N$ sequence where initial states $(h_{0}^{stm},h_{0}^{ltm})$ are learnable parameters.
@@
To provide the Generator estimates of the active state realizations we sample $\mathbf{\hat{r}}$ from the Classifier’s final belief-state $B_{T}$. We assume that the Classifier has already acted and insert the taken action $a$ into the $g$’th index of $\hat{\mathbf{r}}$. We also provide the Classifier’s confidence in the sampled realizations $\hat{\mathbf{r}}_{conf}=B_{T\hat{\mathbf{r}}}$. Depending on whether $a$ was rewarded or not, the $g$’th index of $\hat{\mathbf{r}}_{conf}$ is set to $1$ or 0.
@@ A.8 Pseudo-code
Input: Total States $|\mathcal{S}|$, Context size $C$, Realizations $R$, Time Horizon $T$
@@
for active state pair $c,c^{\prime}\in C$ do

### 20 Apr 2026 v3

@@ A.4 Dis-entanglement
Crucially, because different realizations of the same variable are mutually exclusive, the likelihood of a specific $r=r^{\prime}$ is perfectly isolated from all other potential realizations and $P_{Z}(r=r^{\prime}\mid\mathbf{o})$ may be updated independently. Next, consider a context of two latent variables ($C=2$). Here the interactions $Z^{i}=(z^{i}_{gc},z^{i}_{cg})$ map to a joint space via the bilinear tensor product $f(z^{i}_{gc})\otimes f(z^{i}_{cg})\in\mathbb{R}^{R\times R}$, forming an $R\times R$ matrix of belief updates. The bilinear construction generates synergy. In other words, the combination of variables loses predictive power if the variables are considered in isolation, i.e. the whole is greater than the sum of its parts [74, 6]. We can see this because evaluating the likelihood of $o^{i}=1$ given a realization $r_{g}=r^{\prime}$ now involves projecting indicator $\mathbf{e}_{r_{c}}$ onto a basis shared with $\mathbf{e}_{r_{g}}$:
$P_{Z^{i}}(o^{i}=1\mid r_{g}=r^{\prime},r_{c})=\mathbf{e}_{r^{\prime}}^{T}\sum_{r_{c}}\left(f(z^{i}_{gc})f(z^{i}_{cg})^{T}\right)\mathbf{e}_{r_{c}}.$
In summary, the entanglement between latent variables emerges from projecting onto a shared basis.
Next we turn to investigating entanglement in real-time. Since SII is an accumulating measure (best used for comparing full trajectories) it is not well suited for high temporal fidelity analysis. Instead, we aim to compare the Naive marginal likelihood $p_{Z}(\mathbf{o}_{t}\mid r_{g})$ and the history-conditioned marginal likelihood $p_{Z}(\mathbf{o}_{t}\mid r_{g},\mathbf{o}_{1:t-1})$. The history-conditioned marginal likelihood can be approximated empirically using the dynamics of the Optimal Observer’s marginalized beliefs:
$p_{Z}(\mathbf{o}_{t}\mid r_{g},\mathbf{o}_{1:t-1})=\sum_{r_{c}}p_{Z}(\mathbf{o}_{t}\mid r_{g},r_{c})p(r_{c}\mid r_{g},\mathbf{o}_{1:t-1})\approx\frac{B_{t+1}(r_{g})}{B_{t}(r_{g})}.$
Here we point out that the Optimal Bayesian update is equivalent to the Naive update ($\frac{B_{t+1}(r_{g})}{B_{t}(r_{g})}\propto p_{Z}(\mathbf{o}_{t}\mid r_{g})$) when the belief dynamics are Markovian with respect to the marginals, meaning that the transition from $B_{t}(r_{g})$ to $B_{t+1}(r_{g})$ is independent of $\mathbf{o}_{1:t-1}$.
The history-dependence of the Optimal Bayesian update can be attributed to the prior’s impact on the functional likelihood, $\sum_{r_{c}}p_{Z}(\mathbf{o}_{t}\mid r_{g},r_{c})p(r_{c}\mid r_{g},\mathbf{o}_{1:t-1})$, manifesting as entanglement between variables $r_{g}$ and $r_{c}$. To quantify this ”non-Markovian-ness”, we measure the Jeffreys Divergence, $J(p,q)={D}_{\mathrm{KL}}(p\;\|\;q)+{D}_{\mathrm{KL}}(q\;\|\;p)$, between (normalized) Naive marginal likelihoods and history-conditioned marginal likelihoods, which we refer to as dis-entanglement:
$\text{Dis-entanglement}_{t}=J\bigl(\frac{p_{Z}(\mathbf{o}_{t}\mid r_{g})}{z_{a}}\;||\;\frac{p_{Z}(\mathbf{o}_{t}\mid r_{g},\mathbf{o}_{1:t-1})}{z_{b}}\bigr)\approx\!J\bigl(\frac{p_{Z}(\mathbf{o}_{t}\mid r_{g})}{z_{a}}\;||\;\frac{B_{t+1}(r_{g})}{B_{t}(r_{g})z_{b}}\bigr)\!,$
where $z_{a}$ and $z_{b}$ are normalization factors to ensure valid probabilities.
To investigate the response of the network dynamics to dis-entanglement events, we performed a cross correlation between dis-entanglement of Optimal Bayesian beliefs and several properties of the network output.Figure 14a demonstrates that dis-entanglement is associated with a change of activity in population space, in both Fully Trained and Echo State networks. However, the magnitude of the response differs greatly between networks (Figure 14b). Dis-entanglement is associated with an increase in activity in the Fully Trained network, in contrast to Echo State activity which appears to shrink. Furthermore, Figure 14c demonstrates that the Fully Trained network’s increase in activity is distributed throughout the population, rather than along a few directions. The combined effect can be observed in Figure 14c, which shows the evolving Participation Ratio of the recurrent activity, which exhibits a spike in dimensionality for the Fully Trained network. This dynamic spread of information could account for the increased RNN dimensionality observed in Figure 11d. These results imply that active control of dimensionality may be a mechanism for disentanglement.
Figure 14: Dis-entanglement correlates with learned dynamics of dimensionality. Cross-correlations between Bayesian dis-entanglement and network (a) absolute distance, (b) L2 norm, (c) Hoyer’s sparsity and (d) Participation Ratio of network dynamics. Only data from $20\leq t\leq T$ was analyzed to isolate the steady-state response profile.
@@ A.5 Representation Classification Chains and Thousand Brains Theory
The Thousand Brains Theory of intelligence [33] postulates that individual cortical columns autonomously model through compositions of compositions. Tasked specifically with learning useful relationships, Representation Classification Chains can be viewed as the potential microcircuit of such a system. This architecture-agnostic circuit generates abstractions $\mathbf{r}$, to explain inputs $\mathbf{o}$, in a manner that enables optimal control of inputs given their internal value, without needing to interact with the environment. These chains incorporate the relational information of embeddings, decision-making of Classifiers, latent representation of generative models and embodiment of a Controller. Importantly, these modules are not combined arbitrarily, but have a principled grounding in the task of compositional generalization to novel combinations of latent variables.
Figure 15: Additional learning trajectories of the Controller trained Offline w/ Generator. Representative examples of the Controller exploring an internally generated Cognitive Gridworld.
Figure 16: Network results extend to environments with 3 interacting variables. (a) Same as Figure 3b-d, for $1,2$ and 33 relevant variables. (b) Divergence of network marginal beliefs from Bayesian marginal beliefs. Over the course of inference, the Fully Trained network’s beliefs (solid lines) diverge from Naive Bayes (on the x-axis) while staying aligned with Joint Bayes (on the y-axis). Conversely, the Echo State network’s beliefs (dashed lines) exhibit the opposite trend, aligning closer to Naive Bayes beliefs. Line colors correspond to contextual variables $C=2$ and $C=3$.
@@ A.6 Environment Hyperparameters
  • $|\mathcal{S}|$ (Total latent variables / states): 500
@@ A.7 Additional Model Hyperparameters
  • Hidden dimensionality of all networks ($N$): 1000
@@
  • RCC Entropy bonus ($\beta_{1}$): $0.1$
@@
  • RCC Entropy bonus ($\beta_{1}$): $0.01$
  • Controller Learning rate: 0.005
  • Controller Training episodes: 2000
@@
  • Controller Entropy bonus ($\beta_{2}$): $0.05$
  • Online Generator Learning Rate: 0.001
@@
When compositional embeddings are learned, training reliability was improved when the learnable embedding estimates were over-parameterized. The relevant variable embedding estimates were passed through learnable linear layers $f_{K}(\cdot),f_{Q}(\cdot)\in\mathbb{R}^{d_{\hat{\mathcal{E}}}\times d_{\mathcal{E}}}$, to project them down from $d_{\hat{\mathcal{E}}}=N$ to dimensionality $d_{\mathcal{E}}$. Estimated interactions were computed from the projections:
@@
$\mathcal{L}_{\text{reg}}=\frac{1}{C\times d_{o}}\sum_{c,i}(\|f_{K}(\hat{\mathbf{k}}^{i}_{c})\|_{2}-1)^{2}+(\|f_{Q}(\hat{\mathbf{q}}_{c}^{i})||_{2}-1)^{2}$
@@
    LSTM: $h_{t}^{stm},h_{t}^{ltm}=\text{LSTM}(h_{t}^{in},(h_{t-1}^{stm},h_{t-1}^{ltm}))$ processes the $T\times N$ sequence where initial hidden states $(h_{0}^{stm},h_{0}^{ltm})$ are learnable parameters.
@@
To provide the Generator estimates of the relevant variable realizations we sample $\mathbf{\hat{r}}$ from the Classifier’s final belief-state $B_{T}$. We assume that the Classifier has already acted and insert the taken action $a$ into the $g$’th index of $\hat{\mathbf{r}}$. We also provide the Classifier’s confidence in the sampled realizations $\hat{\mathbf{r}}_{conf}=B_{T\hat{\mathbf{r}}}$. Depending on whether $a$ was rewarded or not, the $g$’th index of $\hat{\mathbf{r}}_{conf}$ is set to $1$ or 0.
@@ A.8 Pseudo-code
Input: Total variables / states $|\mathcal{S}|$, Context size $C$, Realizations $R$, Time Horizon $T$
@@
for relevant variable pair $c,c^{\prime}\in C$ do